The measure of an angle is seventy-four times the measure of a supplementary angle. What is the measure of each angle?

Answer by algebrahouse.com(1659)

x = measure of the angle
180 - x = measure of supplement {supplementary angles add up to 180°}

x = 74(180 - x) {measure of the angle is 174 times the supplement}
x = 13,320 - 74x {used distributive property}
75x = 13,320 {added 74x to each side}
x = 177.6 {divided each side by 177.6}

the angle is 177.6°
the supplement is 2.4°

For more help from me, visit: www.algebrahouse.com